OpenStudy (anonymous):

I have a solution to an algebra problem, but it is in the form of a decimal (0.32) and (2.92), the problem is, is that my teacher want the answer to be integers or fractions, how can I put those two answers in the form of integers or a fractions?

OpenStudy (anonymous):

0.32 = 32/100 = 8/25 2.92 = 2 92/100 = 2 23/25

OpenStudy (anonymous):

happy with that ?

OpenStudy (anonymous):

Hi Jimmy, it looks good, but I am not sure if we can use mixed number, but if that is the only way then that is what I will use. Thanks a lot. Your are Great!!!!

OpenStudy (anonymous):

no probs- u can convert the mixed number to improper if u like 2 23/25 = 73/25

OpenStudy (anonymous):

good luck!